### v2

# Determination of pairing matrix elements from average single particle level densities

Abstract
A simple and efficient method to treat nuclear pairing correlations within a simple Hartree-Fock–plus-BCS description is proposed and discussed. It relies on the fact that the intensity of pairing correlations depends crucially on level densities around the Fermi surface ($\rho(e_{F})$) and that any fitting of nuclear energies as functions of the nucleon numbers is akin of a semi-classical average, smoothing out their quantal structure. A particular attention has been paid to two points generally ignored in previous similar approaches. One is a correction advocated by Möller and Nix [Nucl. Phys. A 536, 20 (1992)] taking into account the fact that the data included into the fit correspond to $\rho(e_{F})$ values systematically lower than average. The second is due to a systematic overestimation of the proton sp level density at the Fermi surface resulting from the local Slater approximation of the Coulomb exchange terms in use in most microscopic descriptions. Our approach is validated by the agreement with data of corresponding calculated moments of inertia of well and rigidly deformed rare-earth nuclei, evaluated according to the Inglis-Belyaev ansatz with some crude Thouless-Valatin corrections. Indeed, the agreement which is found, is at least of the same quality as what results from a specific fit of the pairing intensities to these particular pieces of data. While this approach is currently limited to the very simple seniority force pairing treatment, it may serve as a starting point to define pairing residual interactions from averaged odd-even mass differences data, using merely average sp level densities associated to calculated canonical basis.
I Introduction
We aim at discussing a simple and phenomenologically successful approach to determine the intensity of a pairing residual interaction used in a two-steps self-consistent approach of low energy nuclear structure as currently performed and sketched below.
A self-consistent mean field is calculated upon using any particle-hole effective interaction (here the Skyrme parametrisation will be considered for the strong interaction part). This defines a single particle (sp) canonical basis from which pairing correlations are introduced either within a self-consistent Hartree-Fock (HF) approach (see the seminal paper [1]) or within a diagonalisation in a highly truncated particle-hole basis (the so-called HTDA as introduced in [2]).
In order to do so, one makes an appropriate choice of the residual interaction: seniority force (constant pairing matrix elements around the chemical potential), delta interaction possibly with a density dependence to enhance the surface effects [3], Gaussian in $r$-space or separable in $p$-space [4]. Here, as a first step, to demonstrate the validity and performance of the general approach we will consider the simple seniority force ansatz.
Bohr, Mottelson and Pines [5] have pointed out, in particular, two nuclear spectroscopic properties strongly contingent upon pairing correlations (odd-even mass differences $\delta E$ and moments of inertia $\mathcal{J}$ (MoI), i.e., in practice, the first 2+ energies in well and rigidly deformed nuclei). Both are a priori accessible to a theoretical description within the above defined self-consistent approaches.
In a previous paper [6], making two independent fits of these two properties in the rare earth region, it has been shown that one has obtained very similar values for the parameters of the seniority force matrix elements. This provides hints that: a) as expected, pairing correlations are indeed (all things being kept similar) the main factor to yield precise values of $\delta E$ as well as $\mathcal{J}$, and b) that such a simplified approach, also followed here, was suited globally to describe these properties.
However, in both approaches the quality of the agreement with data was locally (i.e. around a given nucleus) contingent upon a perfect reproduction of the ordering and fine distribution of the sp energies. But a fit corresponds in its very principle to a reproduction of some properties on the average. The above quoted success was obtained upon considering reasonably sized samples so one could conclude that on the average the regime of pairing correlations was correctly adjusted.
Taking stock of this remark, we have chosen here to perform an estimate of the pairing strengths using a tool a priori less sensitive to local sp spectrum deficiencies. The question we ask ourselves here is whether in doing so, we would obtain a reproduction of MoI at least as good as what has been obtained by a direct fit of these moments (see Ref. [6]).
In practice we used the so-called uniform gap method ([7], see also Ref. [8]) providing a value of the average matrix elements (around the Fermi energy) of the pairing residual interaction upon adopting an adequate smooth parametrisation of the neutron and proton pairing gaps $\Delta(N,Z)$.
Some special attention has been paid to two important points of a different nature.
The first issue is of a physical nature. Some analytical forms of the $\Delta(N,Z)$ function have been proposed by Jensen et al. [9] and Madland and Nix [10], but Möller and Nix [11] have pointed out a systematic bias of these approaches related with the lower than average character of the sp level density at equilibrium deformations. They proposed a phenomenological correction which we will adopt in this paper.
The second point concerns an approximation made in most self-consistent calculations (including most of ours) for the sake of numerical easiness. In previous studies (e.g. in Table II of [6]) one has noticed that the treatment of pairing correlations was significantly more successful for neutrons than it was for protons. It has been hinted that it was related to an approximate treatment of the non-local Fock term of the Coulomb mean field due to Slater [12]. As discussed in Sub-Section II.4, this approximation systematically overestimates the sp level density near the Fermi level [13, 14, 15]. A proper account of this spurious enhancement was thus in order and an appropriate correction to the proton sp level density has been implemented upon comparing the effect of the Coulomb exchange terms issued from approximate and exact calculations.
Calculations have been performed for a sample of $19$ lanthanide nuclei supplemented by three isotopes of Hafnium and one isotope of Tungsten (hereafter loosely refered to as rare earth nuclei) and 22 actinide nuclei. They are listed in Table 1. These nuclei are well ($\beta_{20}$ values in the $0.2-0.3$ range with usual notation) and rigidly deformed. The latter property is ascertained by a ratio of their first $2^{+}$ and $4^{+}$ levels $R_{42}\geq 3.290$ as displayed in Table 1 (energy data taken from the compilation of Ref. [16]).
Table 1: The ratio of the first $4^{+}$ over $2^{+}$ energy $R_{42}$, the estimated neutron $V_{n}$ and proton $V_{p}$ pairing matrix elements (in MeV), the calculated MoI (in $\hbar^{2}/\mbox{MeV}$), including a Thouless-Valatin corrective factor of $\alpha=1.32$, obtained with the SIII, SkM* and SLy4 parametrisations and the experimental MoI extracted from the first $2^{+}$ energy, $\mathcal{J}_{exp}$, for a nucleus with $Z$ protons and $N$ neutrons. Actinide nuclei removed from the MoI r.m.s analyses presented in Table 5 are marked with dashed lines.
Z N A $R_{42}$ SIII SkM* SLy4 $\mathcal{J}_{exp}$
$V_{n}$ $V_{p}$ $\mathcal{J}_{TV}$ $V_{n}$ $V_{p}$ $\mathcal{J}_{TV}$ $V_{n}$ $V_{p}$ $\mathcal{J}_{TV}$
62 94 156 3.290 0.1901 0.2529 37.742 0.1769 0.2450 39.438 0.2013 0.2537 38.221 39.531
62 96 158 3.301 0.1862 0.2517 38.068 0.1727 0.2440 42.063 0.1964 0.2527 37.148 41.209
62 98 160 3.292 0.1818 0.2510 40.765 0.1683 0.2435 40.693 0.1916 0.2518 39.215 42.373
64 96 160 3.302 0.1865 0.2454 36.537 0.1735 0.2383 39.943 0.1970 0.2467 35.192 39.860
64 98 162 3.302 0.1822 0.2442 40.663 0.1694 0.2377 38.840 0.1925 0.2453 38.258 41.899
64 100 164 3.295 0.1776 0.2430 40.657 0.1651 0.2370 38.136 0.1880 0.2445 42.849 40.944
64 102 166 3.297 0.1738 0.2422 42.775 0.1612 0.2366 39.549 0.1843 0.2437 37.230 42.857
66 96 162 3.294 0.1867 0.2389 35.283 0.1742 0.2308 38.742 0.1974 0.2393 35.302 37.193
66 98 164 3.301 0.1825 0.2379 39.453 0.1701 0.2302 38.577 0.1931 0.2381 39.004 40.876
66 100 166 3.310 0.1783 0.2373 38.850 0.1645 0.2294 38.936 0.1888 0.2372 43.868 39.171
66 102 168 3.313 0.1743 0.2367 40.901 0.1564 0.2290 42.649 0.1847 0.2359 38.845 40.021
68 100 168 3.309 0.1797 0.2318 36.297 0.1669 0.2224 37.682 0.1893 0.2319 41.603 37.592
68 102 170 3.310 0.1749 0.2307 39.076 0.1631 0.2217 39.339 0.1854 0.2302 37.400 38.173
68 104 172 3.314 0.1712 0.2296 35.341 0.1605 0.2209 39.610 0.1796 0.2287 36.688 38.961
70 100 170 3.293 0.1791 0.2261 35.754 0.1674 0.2167 38.655 0.1895 0.2266 40.594 35.606
70 102 172 3.305 0.1753 0.2246 37.865 0.1637 0.2154 41.260 0.1859 0.2244 37.624 38.099
70 104 174 3.310 0.1717 0.2227 35.636 0.1602 0.2135 43.698 0.1823 0.2221 37.465 39.231
70 106 176 3.310 0.1681 0.2215 34.882 0.1567 0.2120 42.205 0.1786 0.2202 38.821 36.525
70 108 178 3.310 0.1646 0.2204 37.397 0.1533 0.2109 42.128 0.1756 0.2185 38.506 35.714
72 106 178 3.291 0.1686 0.2164 31.988 0.1573 0.2108 33.572 0.1789 0.2188 33.180 32.196
72 108 180 3.307 0.1652 0.2154 33.615 0.1539 0.2097 31.557 0.1753 0.2175 31.464 32.146
72 110 182 3.295 0.1617 0.2144 31.436 0.1506 0.2089 31.185 0.1725 0.2165 31.184 30.678
74 108 182 3.291 0.1653 0.2112 30.384 0.1543 0.2061 27.708 0.1753 0.2141 27.755 29.968
90 144 234 3.291 0.1216 0.1744 57.708 0.1131 0.1671 65.546 0.1297 0.1715 74.584 60.545
92 140 232 3.291 0.1254 0.1632 63.879 0.1172 0.1628 - 0.1340 0.1678 76.273 63.061
92 142 234 3.296 0.1235 0.1698 - 0.1153 0.1630 - 0.1291 0.1677 82.008 68.969
92 144 236 3.304 0.1217 0.1700 63.052 0.1135 0.1634 - 0.1300 0.1678 84.126 66.307
92 146 238 3.303 0.1203 0.1701 59.663 0.1117 0.1639 - 0.1281 0.1682 71.416 66.791
92 148 240 3.347 0.1181 0.1704 - 0.1099 0.1642 - 0.1262 0.1691 63.613 66.667
94 142 236 3.304 0.1237 0.1654 - 0.1156 0.1599 71.615 0.1291 0.1593 - 67.219
94 144 238 3.312 0.1219 0.1654 67.935 0.1138 0.1601 70.205 0.1250 0.1645 - 68.081
94 146 240 3.309 0.1201 0.1656 65.827 0.1120 0.1602 73.433 0.1284 0.1647 71.278 70.054
94 148 242 3.307 0.1183 0.1703 61.921 0.1103 0.1606 - 0.1265 0.1650 66.699 67.355
94 150 244 3.391 0.1167 0.1663 63.423 0.1086 0.1609 64.584 0.1248 0.1654 73.284 67.873
94 152 246 3.308 0.1150 0.1669 - 0.1075 0.1612 59.134 0.1233 0.1656 59.224 64.240
96 146 242 3.252 0.1202 0.1622 - 0.1123 0.1565 - 0.1290 0.1611 - 71.208
96 148 244 3.314 0.1185 0.1626 66.225 0.1061 0.1566 - 0.1267 0.1610 70.430 69.837
96 150 246 3.313 0.1169 0.1630 65.332 0.1091 0.1569 68.096 0.1250 0.1612 - 70.008
96 152 248 3.313 0.1152 0.1635 - 0.1073 0.1571 64.122 0.1197 0.1614 65.674 69.124
98 150 248 3.318 0.1170 0.1597 63.764 0.1093 0.1532 71.893 0.1251 0.1592 - 72.237
98 152 250 3.321 0.1153 0.1599 - 0.1076 0.1534 68.774 0.1234 0.1578 67.810 70.223
98 154 252 3.319 0.1138 0.1640 - 0.1060 0.1537 62.400 0.1216 0.1579 61.026 65.617
100 154 254 3.319 0.1139 0.1560 - 0.1063 0.1499 64.823 0.1218 0.1546 62.273 66.679
100 156 256 3.317 0.1098 0.1560 - 0.1047 0.1502 59.527 0.1201 0.1545 61.250 62.344
102 150 252 3.310 0.1152 0.1520 67.468 0.1097 0.1469 - 0.1252 0.1521 - 64.655
The paper is organised as follows. The general fitting approach is presented in Section II. The extraction of averaged sp level density is performed à la Strutinsky [7] from our microscopic calculations for a sample of well and rigidly deformed nuclei. The uniform gap method is used to extract the matrix element of the pairing residual interaction upon using the Möller-Nix ansatz for average $\Delta(N,Z)$ values. The definition of effective average gaps to be fitted as well as the correction in the proton case for the approximation made on the Fock Coulomb terms are also discussed there. Some technical details are briefly presented in Section III. They include the choice of the sample of deformed nuclei in the two considered regions of heavy deformed nuclei (around rare-earth and actinide elements) and the specific choice made for the Strutinsky averaging of the sp level density. Our results obtained with three parametrisations of the Skyrme interaction, namely SIII [17], SkM* [18] and SLy4 [19], for the moments of inertia are presented in Section IV. Finally, Section V summarizes the main conclusions of our work.
II The approach
II.1 Overview of the approach
Assuming that we know the smooth behaviour of the average neutron and proton pairing gaps with $N$ and $Z$, we devise here an approach to get the corresponding smooth evolution of pairing matrix elements $V_{q}$ (where $q$ stands for the charge states) averaged over a given sp valence space being contained in the $[\lambda_{q}-\Omega,\lambda_{q}+\Omega]$ range where $\lambda_{q}$ is the Fermi energy to be defined later, while the above energy interval (spanning a $2\Omega$ energy range) characterizes the domain of sp states (of the canonical basis) active in the BCS treatment.
From the exact sp level density $\rho(e)$ as a function (rigorously speaking distribution) of the energy $e$, we define a semi-classical sp level density function $\tilde{\rho}(e)$ obtained in practice by a Strutinsky averaging in $e$ [7]. We recall here the close connection of the Strutinsky energy averaging with a semi-classical averaging à la Wigner-Kirkwood (see Ref. [20]).
As stated in the introduction, we restrict ourselves in this work to constant pairing matrix elements $V_{q}$ for each charge state $q$ within an energy range of $2\Omega$ centered around averaged Fermi energies $\tilde{\lambda}_{q}$ defined below. Limiting the interval of sp states active in the BCS variational process makes the value of $V_{q}$ dependent on the value of $\Omega$, as well known. Here consistently we will take $\Omega=6$ MeV.
The matrix elements $V_{q}$ are determined in terms of the averaged sp level densities $\tilde{\rho}_{q}(e)$ and the average pairing gaps $\tilde{\Delta}_{q}$ through the following gap equation:
[…]
The Fermi energies $\tilde{\lambda}_{q}$ are defined from the average density ${\tilde{\rho}}_{q}(e)$ for a total fermion number $N_{q}$ such that
[…]
A frequently used approximation of the above gap equation consists in assuming that the variation of ${\tilde{\rho}}_{q}(e)$ is small enough within the $[\tilde{\lambda}_{q}-\Omega,\tilde{\lambda}_{q}+\Omega]$ sp energy interval so that one can replace it by a constant, namely its Fermi energy value ${\tilde{\rho}}_{q}(\lambda_{q})$. Upon performing the integral one then obtains a closed form formula for $V_{q}$ as
$\frac{1}{V_{q}}=\frac{1}{2}{\tilde{\rho}}_{q}(\lambda_{q})sinh^{-1}\Big{(}\frac{\Omega}{{\tilde{\Delta}}}_{q}\Big{)}.$ (3)
This approximation is shown to be indeed rather good as displayed in Figure 1 where we have plotted the difference in the pairing matrix elements obtained using equation (1) and (3) using an integration interval defined by $\Omega=6$ MeV. The difference in the proton pairing matrix elements between the two equations is mostly localized between ±$\pm 1$ keV while the largest difference for neutrons in absolute value is $4.3$ keV.
While we have shown en passant that the approximate equation (3) is a rather good approximation, we have however resorted to making a full integration using equation (1) for our calculations.
At the end of this process, we will then have for each charge state an average matrix element of the pairing residual interaction $V_{q}(N,Z)$ as a function of $N$ and $Z$.
Figure 1: Difference in pairing matrix elements $\Delta v_{q}$ between values obtained using a full integration over the pairing window $\Omega$ and those obtained using the asinh function given in equation (3)
II.2 Determination of the average single-particle level density
The crucial part is then to determine the average level density for a given set of sp levels. We have computed the average level density using the equation [7, 21]
[…]
The so-called curvature corrections (as discussed in Refs. [7, 21]) are taken care of by the $f(x)$ term defined as
$f(x)=P(x)\>w(x)$ (5)
where $P(x)$ is a polynomial of degree $2M$ in $x$ defined in terms of generalized Laguerre polynomials $L_{M}^{(\alpha)}$ of the form
$P(x)=L_{M}^{1/2}(x^{2})=\sum_{n=0}^{M}a_{2n}\>x^{2n}.$ (6)
with coefficients $a_{2n}$ given in Table 2 and $w(x)$ being a weightage of Gaussian type defined by
$w(x)=\frac{1}{\sqrt{\pi}}e^{-x^{2}}.$ (7)
M $a_{0}$ $a_{2}$ $a_{4}$ $a_{6}$ $a_{8}$
0 1
1 3/2 -1
2 15/8 -5/2 1/2
3 35/16 -35/8 7/4 -1/6
4 945/384 -315/48 63/16 -9/12 1/24
Table 2: Five lowest generalized Laguerre polynomial coefficients entering equation (6).
The value of the smoothing width $\gamma$ appearing in equation (4) is crucial to define a correct energy window for the discrete sp levels to be considered in the integration of equation (4). This, as discussed e.g. in Refs. [7, 21], is generally defined by fulfilling the so-called plateau condition ensuring that the shell effect energy is almost constant as a function of $\gamma$.
In practice, we chose a value of $\gamma$ (in MeV) according to the total nucleon number $A$ as
$\gamma=\beta\>\frac{41}{A^{1/3}}[MeV]$ (8)
using the standard (see Ref. [22]) formula for the energy spacing between major oscillator shells. The specific value of the above constant $\beta$ and the polynomial order $2M$ in $x$ of $P(x)$ will be discussed in Sub-Section III.2.
II.3 Empirical vs effective-pairing gaps
The next crucial ingredients to the approach are the pairing gaps extracted from the data on $\delta E$, entering equation (1) or (3). At this point, it is important to differentiate between empirical pairing gaps as used e.g. in Refs. [9, 10] and those to be employed in any fitting approach such as ours. The latter are referred to as effective interaction pairing gaps in [11].
As discussed in [11], these effective gap values should take into account a bias due to the shell effects in the sp level density at equilibrium deformation making it systematically lower than its average value. Therefore the experimental pairing gaps should not be used as such to determine average pairing properties. Their values should be quenched since upon not doing that a fit of the residual interaction on them would lead to an overestimation of the pairing correlations in actual Hartree-Fock–plus–BCS (HF+BCS) calculations.
Herein, we use the effective-interaction average pairing gaps, phenomenologically determined in Ref. [11] to be:
$\tilde{\Delta}_{q}=\frac{rB_{s}}{N_{q}^{1/3}}$ (9)
where $B_{s}$ is set to $1$ and $r=4.8$ MeV [11].
II.4 Effective pairing gaps in the proton case
In the case of protons, the above value of $r$ yields slightly too high BCS pairing gaps since the above formula corresponds supposedly to an exact treatment of the Coulomb interaction This is, however, not the case in most mean-field calculations. In order to avoid considering non-local mean fields, the Coulomb-exchange contribution is usually accounted for à la Slater, i.e. upon using the infinite nuclear matter Pauli correlation function [12]. As found long ago [13] and confirmed later [14, 15] the Slater approximation in use, systematically overestimates the sp level density near the Fermi level with respect to the corresponding exact treatment.
In order to correct for this systematic spurious trend modifying significantly the level density, we have proceeded as follows. We have considered for a given nucleus, two sp proton spectra obtained within the Skyrme HF+BCS framework, corresponding to an exact treatment and the Slater approximation of the Coulomb exchange terms. The exact calculation of the Coulomb energy matrix elements has been performed according to the method developed in Refs. [13] and [15]. This has yielded a corresponding ratio $R_{p}$ of the BCS gaps:
$R_{p}=\frac{\Delta_{p}^{exact}}{\Delta_{p}^{Slater}}.$ (10)
We take for granted the average effective gap of Ref. [11] (see equation (47)) with $r=4.8$ MeV as reproducing adequately effective average pairing matrix elements. For a given well and rigidly deformed nucleus and a given interaction, we perform calculations with the Slater approximation using the Möller-Nix value of $r$ and reasonable values (as defined below) of the average pairing matrix element for instance as (see the discussion of such a choice in Subsection III.1)
It is well known that the sp spectra are almost unaffected by a variation of $V_{q}^{0}$ in quite a large range of values of around values such as those displayed above in eq. (11). This is demonstrated at least for the average sp level densities at the Fermi surface in Table 3. For two nuclei in their deformed ground states, 176Yb and 240Pu, we have calculated $\tilde{\rho}_{q}(\lambda_{q})$ for three values of the intensity parameter $G_{q}=16,19,22$ MeV. They vary for both nuclei and for all values of $G_{q}$ no more than $0.1\%$. Therefore, the resulting values of the average pairing matrix elements $V_{q}^{1}$ contingent merely upon the sp spectra should seemingly not depend on the arbitrary chosen values of $V_{q}^{0}$. This is clearly so for the neutron matrix elements $V_{n}$ and not really the case for the proton matrix elements $V_{p}$, for reasons which will be discussed now.
$V_{q}^{0}=\frac{G_{q}}{11+N_{q}}\;\;\;\mbox{with}\;\forall_{q},\;G_{q}=19\;\mbox{MeV}.$ (11)
Table 3: The average Fermi level densities for neutrons and protons (in MeV-1) are reported in columns 4 and 5, respectively, for the 176Yb and 240Pu nuclei in the ground-state deformation characterized by its quadrupole moment $Q_{20}$ (in barns). Columns 6 and 7 give the corresponding fitted values (in MeV) of the average matrix elements $V_{q}$ (with $M=2$, $\beta=1.2$). These values are generated from HF+BCS calculations using different $G_{q}$ values (equal for neutrons and protons) as reported in column 2.
Nucleus $G_{q}$ $Q_{20}$ ${\tilde{\rho}}_{n}(\lambda_{n})$ ${\tilde{\rho}}_{p}(\lambda_{p})$ $V_{n}$ $V_{p}$
Yb-176 16 19.13 4.709 3.619 0.1681 0.2151
19 18.73 4.710 3.622 0.1681 0.2215
22 18.34 4.717 3.625 0.1680 0.2300
Pu-240 16 28.72 6.388 4.736 0.1201 0.1602
19 28.26 6.386 4.735 0.1201 0.1656
22 27.77 6.385 4.735 0.1201 0.1733
It is not as simple as to perform calculations with these $V_{p}^{0}$ values taking exactly into account the Coulomb exchange calculations to get $\Delta_{p}^{exact}$ and then determine $R_{p}$ as defined in eq (10) to correct for the Möller-Nix parameter. It turns out that the value of $R_{p}$ depends on the chosen value of $V_{p}^{0}$ or in effect on the degree of pairing correlations as a monotonically increasing function. This could be expected since the more pairing correlations are present, the closer to a smoothed-out level density distribution one would get, making it more and more closer to the one present in infinite nuclear matter corresponding to the Slater approximation.
To quantify the degree of pairing correlations, we consider the so-called pair condensation energy $E_{cond}^{p}$ for protons (the absolute value of the part of the HF+BCS energy which involves the abnormal density), given by
$E_{cond}^{p}=\frac{\Delta_{p}^{2}}{V_{p}}.$ (12)
To determine the variation of $R_{p}$ as a function of $E_{cond}^{p}$, we consider a sample of 19 nuclei in the case of SIII (17 nuclei for SLy4 and SkM*) with respect to the 45 nuclei listed in Table 1. Specifically, we have excluded nuclei exhibiting large sp energy gaps leading to artifically low BCS pairing gap, in view of the well-known deficiency of the BCS approximation in such weak pairing regimes (see e.g. Ref [23]).
For these nuclei, we performed both exact Coulomb and Slater approximation calculations using the pairing matrix elements listed in Table 4. The obtained ratio $R_{p}$ of the BCS proton gaps from all three Skyrme parametrisations considered herein, are then plotted in Figure 2 as a function of the proton condensation energy $E_{cond}^{p}$.
The ratio $R_{p}$ increases, albeit rather minimally, with $E_{cond}^{p}$ defined in equation (12). A fit of the data to a linear equation yields
$R_{p}=0.0181\>E_{cond}^{p}+0.781$ (13)
where $E_{cond}^{p}$ is given in MeV. This equation allows for an estimation of the reduction factor to the Möller-Nix parameter for any Skyrme parametrisation and at a given initial pairing matrix element $V_{p}^{0}$. Multiplying the initial Möller-Nix parameter $r=4.8$ MeV with the reduction factor, one then obtains the proton pairing gap to be utilized in the estimation of pairing matrix element via equation (1).
Table 4: The ratio $R_{p}$ of BCS proton pairing gap between Slater approximation $\Delta_{p}^{Slater}$ and exact Coulomb exchange $\Delta_{p}^{exact}$ calculations using the initial neutron and proton pairing matrix elements listed in column 4 and 5 respectively. The proton condensation energy defined in equation (12) are given in column 9.
$Z$ $N$ $A$ $V_{n}^{0}$ $V_{p}^{0}$ $\Delta_{p}^{Slater}$ $\Delta_{p}^{exact}$ $R_{p}$ $E_{cond}^{p}$
SIII 62 94 156 0.1902 0.2686 1.0449 1.1884 0.879 5.2579
62 96 158 0.1855 0.2682 0.9930 1.1319 0.877 4.7770
64 96 160 0.1862 0.2600 1.0430 1.1894 0.877 5.4411
66 98 164 0.1824 0.2520 1.0326 1.1612 0.889 5.3512
66 100 166 0.1784 0.2515 1.0061 1.1436 0.880 5.1995
66 102 168 0.1742 0.2511 0.9788 1.1254 0.870 5.0449
68 104 172 0.1712 0.2433 0.9818 1.1108 0.884 5.0721
70 104 174 0.1717 0.2363 0.9235 1.0772 0.857 4.9115
70 106 176 0.1681 0.2358 0.8414 1.0169 0.827 4.3853
72 106 178 0.1625 0.2291 0.9270 1.0700 0.866 4.9968
74 108 182 0.1652 0.2227 0.9229 1.0715 0.861 5.1554
92 142 234 0.1235 0.1765 0.9011 1.0114 0.891 5.7972
92 148 240 0.1182 0.1756 0.9965 1.0954 0.910 6.8322
94 148 242 0.1183 0.1720 0.8597 0.9952 0.864 5.7574
94 150 244 0.1167 0.1718 0.9058 1.0375 0.873 6.2669
96 148 244 0.1185 0.1686 0.7969 0.9611 0.829 5.4776
98 154 252 0.1138 0.1646 0.9361 1.0406 0.900 6.5775
100 154 254 0.1139 0.1615 0.8723 0.9700 0.899 5.8271
100 156 256 0.1124 0.1612 0.8765 0.9806 0.894 5.9661
SLy4 62 94 156 0.2013 0.2688 1.2149 1.1111 0.915 5.4913
62 96 158 0.1964 0.2678 1.2015 1.0971 0.913 5.3897
64 96 160 0.1970 0.2603 1.2320 1.1356 0.922 5.8315
66 98 164 0.1931 0.2523 1.1568 1.0111 0.874 5.3031
66 100 166 0.1889 0.2516 1.1287 0.9786 0.867 5.0636
66 102 168 0.1847 0.2507 1.0978 0.9427 0.859 4.8064
68 104 172 0.1815 0.2432 1.0554 0.9247 0.876 4.5793
72 106 178 0.1789 0.2300 1.1540 1.0573 0.916 5.7891
74 108 182 0.1753 0.2237 1.1912 1.0646 0.894 6.3422
92 148 240 0.1263 0.1742 1.0504 0.9242 0.880 6.3322
94 146 240 0.1283 0.1717 0.9364 0.8173 0.873 5.1070
94 148 242 0.1265 0.1712 0.9784 0.8606 0.880 5.5929
94 150 244 0.1248 0.1707 1.0267 0.9148 0.891 6.1745
96 148 244 0.1267 0.1681 0.8995 0.7800 0.867 4.8136
98 154 252 0.1216 0.1637 0.9353 0.7838 0.838 5.3446
100 154 254 0.1217 0.1609 0.9020 0.8601 0.954 5.0579
100 156 256 0.1201 0.1603 0.9253 0.8174 0.883 5.3397
SkM* 62 94 156 0.1770 0.2574 1.2016 1.0758 0.895 5.6088
62 96 158 0.1726 0.2570 1.1636 1.0319 0.887 5.2677
64 96 160 0.1735 0.2493 1.2251 1.1169 0.912 6.0193
66 98 164 0.1701 0.2414 1.1526 1.0119 0.878 5.5037
66 100 166 0.1661 0.2409 1.1356 0.9898 0.872 5.3521
66 102 168 0.1623 0.2405 1.1239 0.9740 0.867 5.2519
68 104 172 0.1594 0.2331 1.0271 0.9068 0.883 4.5248
72 106 178 0.1572 0.2199 1.1150 1.0251 0.919 5.6524
74 108 182 0.1543 0.2138 1.1472 1.0296 0.897 6.1558
92 148 240 0.1100 0.1675 1.0164 0.8614 0.848 6.1661
94 146 240 0.1120 0.1644 0.9456 0.8330 0.881 5.4395
94 148 242 0.1103 0.1642 0.9824 0.8719 0.888 5.8792
94 150 244 0.1086 0.1639 1.0222 0.9149 0.895 6.3749
96 148 244 0.1106 0.1609 0.9179 0.8042 0.876 5.2362
98 154 252 0.1060 0.1570 0.9236 0.7692 0.833 5.4339
100 154 254 0.1063 0.1539 0.8662 0.7278 0.840 4.8742
100 156 256 0.1047 0.1537 0.9012 0.7656 0.850 5.2850
Having at hand the function $R_{p}(E^{p}_{cond})$, we will proceed as follows. We perform HF+BCS calculations with Coulomb à la Slater with the matrix elements $V_{p}^{0}$. We then obtain BCS pairing gaps $\Delta_{q}$ and in particular the proton gap $\Delta_{p}(V_{p}^{0})$ which combined with $V_{p}^{0}$ will provide us with $E_{cond}^{p}(V_{p}^{0})$ and the associated value $R_{p}^{0}$ of the gap ratio defined in eq. (10). It is clear that this ratio depends on the retained value for the initial proton pairing matrix element $V_{p}^{0}$. The necessity for an iterative determination of the correct ratio $R_{p}$ consistent with a corresponding matrix element $V_{p}$ should, in principle, be advocated.
However as it will be discussed at length in Appendix A, the convergence of this iterative process is indeed very fast. Furthermore, upon making some very limited preliminary studies, it is easy to determine a priori, for a given particle-hole interaction, an interval of initial values of $V_{p}^{0}$, previously dubbed as reasonable, such that the particular choices which are made for those, lead to insignificant corrections. It appears that the above mentioned choice of $G_{n}=G_{p}=19$ MeV is convenient in this respect. We therefore stick to the corresponding initial values of
$r_{corr}=r\>R_{p}^{0}$ (14)
without reiterative procedure. We get the final value $V_{p}^{1}$ of the proton average pairing matrix element as done in the first stage of this calculation for a given nucleus and a given interaction, using now the corrected value $r_{corr}$ of the Möller-Nix parameter for the estimation process.
III Technical details
III.1 Generation of sp levels
We have performed HF+BCS calculations for the deformed nuclear ground-states using three Skyrme parametrisations, namely the SIII [17], SkM* [18] and SLy4 [19], for the strong interaction part of the particle-hole interaction.
The canonical basis is determined upon solving the HF equations resulting from the corresponding energy density functional of the one-body reduced density matrix including self-consistently the BCS occupation probabilities. The eigensolutions of the corresponding one-body Hamiltonian are obtained by projection of their eigenstates onto the eigenstates of an axially symmetrical harmonic oscillator, a choice consistent with the axial and intrinsic parity symmetries imposed onto our solutions.
The size of the deformation-dependent basis corresponds for spherical solutions to 17 major shells (i.e. with $N_{0}=16$ in the notation of [1]). The two parameters defining the size and the ellipsoidal deformation of the harmonic oscillator potential (i.e. $b$ and $q$ respectively in the notation of [1]) are optimised for each nucleus to yield the lowest equilibrium energy. Integrals involving the densities are performed using the Gauss-Hermite and Gauss-Laguerre approximate integration methods with 50 and 16 mesh points, respectively.
Pairing correlations are only considered in the isospin $T=1$ channel, which amounts in practice, for the considered nuclei far enough from the $N=Z$ line, to restrict to neutron-neutron and proton-proton pairing (thus for $|T_{z}|=1$). As already mentioned in Section I we define this residual pairing interaction $\widehat{v}_{res}$ from an average of its BCS matrix elements
$V_{q}^{ij}=\langle{i_{q}\bar{i_{q}}}|\widehat{v}_{res}\Big{(}|{j_{q}\bar{j_{q}}\rangle-|\bar{j}_{q}j_{q}}\rangle\Big{)}$ (15)
for a given pair $(i,j)$ of the canonical basis states.
Now, we make a further phenomenological step considering as in Ref. [24] a specific dependence of $V_{q}$ on the neutron or proton numbers $N_{q}$ in the following form
$V_{q}=\frac{G_{q}}{11+N_{q}},$ (16)
and in what follows $G_{q}$ will be referred to as the pairing strength.
The validity of this parametrisation has been demonstrated by the quality of the description at the same time of odd-even mass differences and of moments of inertia obtained in Ref. [6]. A residual interaction is by definition dependent on the number of fermions (through the dependence of the mean field). Yet, it is worth noting that the above parametrisation does not necessarily represent only such a dependence but also, and may be primarily, in an average fashion, the corresponding dependence of the sp wavefunctions (e.g. through their size or compactness).
Thus our task here is to determine, as sketched in Section II, the two parameters $G_{n}$ and $G_{p}$ (and thus $V_{n}$ and $V_{p}$) for each of the 35 well and rigidly deformed nuclei in the rare earth and actinide regions (see Section IV for details). Our approach depends only on sp spectra (through the Fermi levels $\tilde{\lambda}_{q}$ and the value of the average sp level density at these energies $\tilde{\lambda}_{q}$). To get the sp spectra, we have considered pairing residual matrix elements defined by $G_{n}=G_{p}=19$ MeV, as already mentioned.
Figure 2: Ratio of the proton BCS gap between exact Coulomb and Slater approximation calculations $R_{p}$ as a function of the proton pairing condensation energy $E_{cond}$ (given in MeV).
III.2 Choice of coefficients to determine average level density
Two important ingredients entering the equation (1), apart from the choice of pairing gap $\tilde{\Delta}_{q}$ which has been addressed in Section II.3 are the order $M$ of the generalized Laguerre polynomial and the constant $\beta$.
To determine optimal values of $M$ and $\beta$, we performed fits of pairing matrix elements with $\Omega=6$ MeV for two rare earth (166Dy and 176Yb) and two actinide (240U and 252No) nuclei. Figure 3 shows the variation of neutron (top panels) and proton (bottom panels) pairing matrix elements with $M$ and $\beta$ for these nuclei.
We searched for the values of $(M,\beta)$ pairs where a plateau in the average $V_{q}$ matrix elements is roughly achieved to ensure that they remain almost constant upon varying $\beta$. A value of $\beta<1$ is not sufficient to smooth the shell effect (see Figure 4) where some remnants shell effects are still apparent even for $\beta=1$. On the other hand, one must avoid too large $\beta$ values (e.g. such that $\beta\geq 1.6$) to avoid the dubious contribution of unbound sp states poorly approximated by their projection onto a truncated harmonic oscillator basis.
From the results displayed in Figure 3, we have taken as an optimal choice the following values of the smoothing parameters: $M=2$ and $\beta=1.2$.
Figure 3: Variation of neutron (top panels) and proton (bottom panels) pairing matrix elements showing their evolution as a function of the order $M$ for different values of $\beta$.
Figure 4: Average neutron (top panels) and proton (bottom panels) sp level densities (in MeV-1) using equation (1) for the Dy-166 nucleus with $M=2$ and $\beta=1.0,1.2$ and $1.4$ as a function of sp energies. Red arrows indicated the location of the Fermi level.
IV Results
Using the approach discussed above, we calculated the MoI using the Inglis-Belyaev [25] formula with the estimated pairing strengths unique to each nucleus. To account approximately for the Thouless-Valatin selfconsistency correction (see Ref. [26]) the calculated MoI have been multiplied by a factor $1.32$, as suggested in Ref. [27] and shown in previous calculations (see e.g. Ref. [6]) to provide good estimates of this effect, for the three Skyrme parametrisations. The experimental MoI $\mathcal{J}_{exp}$ are determined from the energies of the first $2^{+}$ excited state in the pure rotor limit (these energies are taken from the compilation of Ref. [16] and tabulated in Table 1).
Prior to comparing our calculated MoI with experimental data, we further eliminate some actinide nuclei which exhibit deficiencies in the sp levels spectra. These nuclei marked with dashed lines in Table 1, all of which are in the heavy nuclei region, shows large energy gap located at incorrect nucleon number. In such nuclei, comparing the calculated MoI with experimental data would not be meaningful as the observed deviation is not due to the estimation procedure proposed herein, but rather due to the underlying mean-field solution, providing locally an inadequate sp level distribution at the Fermi surface.
Indeed, the proposed method which relies on an semi-classical averaging is blind to the existence of a large sp energy gap. As examples to this point, we refer to the 246Pu and 248Cm ($N=152$) isotones listed in Table 1. In Ref. [28] it was reported that an incorrect energy gap was found at $N=152$ with the SIII parametrisation. This deficiency is not propagated to the estimated $V_{n}$ as can be clearly seen when comparing the estimated $V_{n}$ with its neighbouring nuclei.
For comparison between our calculated values with experimental MoI, we look at the root-mean-square (r.m.s) deviation $\chi_{\mathcal{J}}$ such that
$\chi_{\mathcal{J}}=\sqrt{\frac{\sum_{i}^{N}(\mathcal{J}_{TV}^{i}-\mathcal{J}_{exp}^{i})^{2}}{N}}$ (17)
where $N$ is the total number of sample nuclei. The r.m.s deviation have been analysed for 23 nuclei around rare-earth region with all three Skyrme parametrisations and 12 (13) actinide nuclei with SIII and SLy4 (SkM*) separately. The $\chi_{\mathcal{J}}$ are tabulated in the top part of Table 5.
We found that the $\chi_{\mathcal{J}}$ values ranges from $1.7$ to $3.0$ $\hbar^{2}/\mbox{MeV}$ in the rare-earth region and from $3.4$ to $4.6$ $\hbar^{2}/\mbox{MeV}$ in the actinide region. It is interesting to note here that the value of $\chi_{\mathcal{J}}\sim 1.77\>\hbar^{2}/\mbox{MeV}$ for rare earth nuclei only is indeed very close to the r.m.s deviation ($1.75\>\hbar^{2}/\mbox{MeV}$) obtained from average pairing strengths fitted to experimental MoI in Ref. [6].
In comparing the $\chi_{\mathcal{J}}$ values, it appears that the agreement with data is less spectacular in the actinide region as compared to the rare-earth region. However, comparison of r.m.s deviation based on $\chi_{\mathcal{J}}$ defined in equation (17) is not suitable since the number $A$ of nucleons are starkly different affecting thus the values of MoI through their $A^{5/3}$ dependence. To remove the $A$ dependence, we compare instead the weighted r.m.s deviation defined as
$\chi_{\mathcal{J}}^{A}=\sqrt{\frac{\sum_{i}^{N}\Big{(}\frac{\mathcal{J}_{TV}^{i}-\mathcal{J}_{exp}^{i}}{A^{5/3}_{i}}\Big{)}^{2}}{N}}$ (18)
$\chi_{\mathcal{J}}^{exp}=\sqrt{\frac{\sum_{i}^{N}\Big{(}\frac{\mathcal{J}_{TV}^{i}-\mathcal{J}_{exp}^{i}}{\mathcal{J}_{exp}^{i}}\Big{)}^{2}}{N}}.$ (19)
In doing so, we see that the agreement with experimental MoI are indeed better for actinides than rare-earth region for the SkM* and SLy4 parametrisations (see Table 5).
Finally, we show that these values of the weighted r.m.s deviation provides a way to estimate the uncertainty in the calculated MoI uniquely for each nucleus. We define two uncertainty ranges for $\mathcal{J}$ associated to a given nucleus displayed in Table 6 as $\Delta_{\mathcal{J}}^{A}$ (resp. $\Delta_{\mathcal{J}}^{exp}$) by multiplying $\chi_{\mathcal{J}}^{A}$ (resp. $\chi_{\mathcal{J}}^{exp}$ ) by $A^{5/3}$ (resp. by $J_{TV}$).
Table 5: Weighted root-mean-square deviations $\chi_{\mathcal{J}}$ (in units of $\hbar^{2}/\mbox{MeV}$), $10^{4}\>\chi_{\mathcal{J}}^{A}$ (in units of $\hbar^{2}/\mbox{MeV}$) and $10^{2}\>\chi_{\mathcal{J}}^{exp}$ as defined in equations (17), (18) and (19 between calculated and experimentally defined MoI obtained with the three Skryme parametrisations.
SIII SkM* SLy4
$\chi_{\mathcal{J}}$ Rare earth 1.769 2.706 2.940
Actinide 4.582 3.394 3.498
$\chi_{\mathcal{J}}^{A}$ Rare earth 3.515 5.080 5.883
Actinide 4.844 3.588 3.621
$\chi_{\mathcal{J}}^{exp}$ Rare earth 4.548 7.230 7.551
Actinide 6.671 5.168 5.248
Table 6: Uncertainty ranges for the moment of inertia (in units of 2superscriptPlanck-constant-over-$\hbar^{2}$/MeV) of some nuclei $\Delta_{J}^{A}$ and $\Delta_{J}^{exp}$ as defined in the text.
Z N $\mathcal{J}_{TV}$ $\mathcal{J}_{exp}$ $\Delta^{A}_{\mathcal{J}}$ $\Delta^{exp}_{\mathcal{J}}$
62 98 40.765 42.373 1.658 1.854
68 100 36.297 37.592 1.798 1.651
72 106 31.988 32.196 1.980 1.455
92 140 63.879 63.061 4.243 4.261
94 150 63.423 67.873 4.615 4.231
102 150 67.468 64.655 4.870 4.501
V Conclusion
In this paper, we have proposed and discussed a simple and efficient method to treat pairing correlations within a microscopic (non-relativistic) description of the structure of atomic nuclei. It takes stock on the fact that the intensity of pairing correlations depends crucially on level densities around the Fermi surface. It is suited to approaches where one has a good knowledge of the particle-hole interaction (e.g. of the usual Skyrme type) yielding (possibly in a self-consistent manner) the normal density matrix (and its canonical basis). Then one searches for a relevant approach to the abnormal density matrix, typically in a HF+BCS framework.
At present, it is limited to a very simple ansatz, namely using constant pairing matrix elements for sp states located in the vicinity of the Fermi surface, dubbed as the seniority force pairing treatment. Moreover, it is only operative a priori, so far, to describe the ground states of well and rigidly deformed nuclei. As a result, it yields, in a well-defined fashion, the pairing average matrix elements suited for a given nuclear ground state and a given particle-hole interaction. It is furthermore important to recall that, by no means it gives any direct access to a residual interaction since its output includes some average information on the wavefunctions of the states around the Fermi surface, particularly their spatial extensions.
For limited that it is now, it may serve, however, as a basis to determine the ingredients of a more elaborate pairing treatment, i.e. defining univocally a residual pairing interaction (contingent now merely on the choice of the particle – hole interaction) and not some average of its matrix elements. This would allow to build up a BCS treatment which can be used more safely on two counts: for all nuclei and away from their equilibrium deformation. On the one hand, it would include explicitely (and not in the average) information on the structure of the relevant sp wavefunctions. And defining an effective residual Hamiltonian, it could be used in a natural fashion away from the limited region where it has made sense to have it fitted with some piece of data, on another hand. The study of this necessary extension is currently under completion and will be presented in a forthcoming publication. To reach that goal it was therefore necessary to assess, first, the quality of the approach discussed here, for what it has been specifically tailored.
To determine the relevant pairing matrix elements, we used smoothly varying (with nucleon numbers) gaps as in Refs. [9, 10] corrected according to the prescription of Ref. [11]. Experimental odd-even mass difference $\delta E$ are the raw data from where these gaps are extracted. Such energies depend of course on the shell structure which is generally well reproduced by state of the art microscopic calculations, yet allowing in some well localised regions to generate some misplacement or bad rendering of sp gap intensities in some low sp level density regions. This is why the information from existing fits of experimental data in terms of smoothly varying (with respect to the nucleon numbers) has not been directly compared with what results from quantal calculations but with their underlying semi-classical content, determined in an approximate fashion. The mere ingredient of the latter relevant to pairing properties within our pairing model is the average sp level density at the Fermi surface and the nucleon numbers of the considered nucleus.
In doing so two important features, absent so far in fits of the pairing intensity, to the best of our knowledge, have been carefully taken into account. One is the correction advocated by Möller and Nix [11] due to the unescapable selection of data corresponding to sp level densities systematically lower than average. The second is due to a systematic overestimation of the proton sp level density at the Fermi surface resulting from the local Slater approximation of the Coulomb exchange contribution to the total energy [13, 14, 15].
The test of our method has consisted in using the so-determined average pairing matrix elements (with three different Skyrme force parametrisations) to compare within an Inglis Belyaev approach (plus some approximate Thouless-Valatin correction) MoI of about forty well and rigidly deformed rare-earth and actinide nuclei, with what is deduced from the experimental energies of their first $2^{+}$ levels.
These sets of Skyrme parametrisations were chosen to allow us to assess the versatility of our pairing matrix elements estimation approach using different mean-field solutions. It was not our intention here to evaluate the quantal deficiency of the sp spectrum produced using any of the Skyrme parametrisations. In fact, we showed that the pairing matrix elements estimated using the proposed approach herein is not affected by the wrong reproduction of the sp energy gap. Therefore, this deficiency is irrelevant for a test of our approach. Consequently, we had to remove these nuclei in our samples because implementing them in our test would unduly affect the quality assessment of the proposed method.
From the remaining list of good sample nuclei, we found a rather good reproduction of the experimental MoI. This is especially so for the case of the Skyrme SIII parametrisation in the rare earth region where our present estimates give excellent agreement with results obtained from a direct fit to these MoI.
This gives confidence on the relevance of what is proposed here and allows to take stock on it to tackle our more ambitious attempt to define pairing residual interactions from averaged $\delta E$ data, using merely average sp level densities at the Fermi surface of the calculated canonical basis.
Acknowledgements.
This work was finalized during a research visit in France funded by the French Embassy in Malaysia via the Mobility Programmes to Support French – Malaysian Cooperation in Research and Higher Education which M.H.K is grateful for. Gratitude also goes to LP2I Bordeaux and IHPC Strasbourg for the warm hospitality extended to him during the visit. M.H.K would also like to acknowlege Universiti Teknologi Malaysia for its UTMShine grant (grant number Q.J130000.2454.09G96).
Appendix A Treatment of the arbitrariness of the initial proton matrix element $V_{p}^{0}$
Here, we discuss how to circumvent the arbitrariness of the chosen initial pairing matrix element $V_{p}^{0}$. Upon calculating the ground state deformation solutions of our selected nuclei within the HF+BCS approach (using the seniority force model) we get proton pairing condensation energies. They are, as already noted, dependent upon the choice of the average pairing matrix elements in use in these calculations. From them, using equation (14) we should define new values of the proton gaps from which using the uniform gap method and the sp spectra, we could generate new proton pairing matrix elements $V^{1}_{p}$. And we could iterate this process to get a convergence of the pairing matrix elements.
The question now is how much change from $V_{p}^{0}$ could we expect for these new matrix elements $V^{1}_{p}$. To estimate that, we assume that we have chosen reasonable values of the original $V_{p}^{0}$, i.e. quantified, for instance, by a deviation of the resulting proton $E_{cond}$ from the value obtained through a converged solution of this iterative process by no more than ± $\pm\>20\%$. We see from Figure 2 that such an interval for $E_{cond}$ corresponds to an interval for $R_{p}$ (and thus on the proton gaps used in the fit) of ± $\sim\pm\>2\%$. From the approximation (shown to be rather good, see Figure 1) of equation 3 concerning the relation between $V_{p}$ and $\Delta_{p}$, we get readily
$\frac{\delta V_{p}}{V_{p}}=\frac{x}{ln(\sqrt{x+\sqrt{x^{2}+1}})}\frac{1}{\sqrt{x^{2}+1}}\frac{\delta\Delta_{p}}{\Delta_{p}}$ (20)
with $x=\Omega/\Delta_{p}$. Using the following values $\Omega=6$ MeV and $\Delta_{p}=1$ MeV, one obtains an uncertainty on $V_{p}$ within the ± $\pm\>0.8\%$ range.
However, a specific convergence study has been performed for the two deformed nuclei considered in Table 3 (176Yb and 240Pu) for the three values of the proton pairing intensity parameter $G_{p}=16,19,22$ MeV. One sees on Table 7 that the value of the proton matrix element $V_{p}$ is converged at the keV level at the second or third iteration. Similarly, the mass quadrupole moment $Q_{20}$ is converged at the fm2 level already at the third iteration, even sometimes at the second. The other lesson comparing the starting value of $V^{0}_{p}$ and the converged one, is that one may hint that a reasonable range for $V^{0}_{p}$ values would lie in between $G_{p}=16$ and $G_{p}=19$ MeV. This constitutes a very simple preliminary study, for any given specific particle-hole interaction, allowing to define a priori for a global study some value of $V^{0}_{p}$ close at a level of about $1\%$ to what an iterative process would produce.
In view of the rough nature inherent to the averaging character of the relation between $E_{cond}$ and $R_{p}$ we deem that the iterative process sketched above presents no solid practical justification to this paper which aims to illustrate a method for determination of pairing strengths with limited dependence on experimental data.
Table 7: Estimated proton pairing matrix elements $V_{p}$ (in MeV) at the ground-state quadrupole moments $Q_{20}$ (in barns) obtained for 176Yb and 240Pu at corresponding iteration number with different starting initial pairing strength $G_{p}^{0}=16,19,22$ MeV. The estimated $V_{p}$ of the preceding iteration are used as initial values for subsequent HF+BCS calculations.
Nucleus Iteration $G_{p}^{0}=16$ MeV $G_{p}^{0}=19$ MeV $G_{p}^{0}=22$ MeV
$Q_{20}$ $V_{p}$ $Q_{20}$ $V_{p}$ $Q_{20}$ $V_{p}$
Yb-176 1 19.13 0.2151 18.73 0.2215 18.35 0.2300
2 18.89 0.2179 18.80 0.2189 18.74 0.2206
3 18.88 0.2183 18.85 0.2185 18.83 0.2188
4 18.86 0.2184 18.86 0.2185 18.85 0.2185
5 18.86 0.2185 18.86 0.2185 18.86 0.2185
Pu-240 1 28.72 0.1602 28.26 0.1656 27.77 0.1733
2 28.35 0.1614 28.34 0.1624 28.32 0.1639
3 28.35 0.1615 28.35 0.1617 28.35 0.1620
4 28.35 0.1616 28.31 0.1617 28.35 0.1616
5 28.35 0.1616 28.35 0.1616 28.35 0.1616